i took 7 tests.
my median score is 90.
the range of scores is 18.
what are the 7 test scores?

Answer by jim_thompson5910(35256)

You would need more information. All you can really find out are the highest and lowest possible scores.


Lowest Score = Median - (range/2) = 90 - 18/2 = 90 - 9 = 81

So the lowest possible score is 81


Highest Score = Median + (range/2) = 90 + 18/2 = 90 + 9 = 99

So the highest possible score is 99